OpenStudy (anonymous):

At the start of a basketball game, the referee tosses a ball into the air for the jump ball. The equation that models the height of the ball is h(t) = -16t2 + 30t + 8. During what time interval is the height of the ball above 10 feet? (Round your answer to the nearest tenth of a second.) a) Between 0.3 and 2.1 seconds b) Between 1.1 and 1.6 seconds c) Between 0.1 and 1.8 seconds

OpenStudy (ajprincess):

substitute 10 h(t) then u will get \[10=-16t^2+30t+8\] \[16t^2-30t+10-8=0\] \[16t^2-30t+2=0\] Divide the whole eqn by 2. then u will get \[8t^2-15t+1=0\] Nw solve for t using the formula \[x=\frac{-b+/-\sqrt{b^2-4ac}}{2a}\]

OpenStudy (ajprincess):

Tell me wat u get when u solve?

OpenStudy (anonymous):

i got c

OpenStudy (ajprincess):

ya right. i too get the same. Is it clear to u?

OpenStudy (anonymous):

yah thanks :)

OpenStudy (ajprincess):

yw.:) At 0.1sec and at 1.8sec it will be at 10ft height.bt between that time interval it will be at a height greater than 10ft.